OpenStudy (anonymous):

How do i multiply (6-3i)(4+9i) and write my answer in standard form?

OpenStudy (anonymous):

@tiffanyyy31 will show you one way, i will show you another \[(a+bi)(c+di)=(ac-bc)+(ad+bc)i\]

OpenStudy (anonymous):

Take for example (a-bi)(a+bi) Using the FOIL method,  Same goes for this equation  6(4)+6(9i)-3i(4)-3i(9i) multiply and combine like terms 24+54i-12i-27i^2 Remember that i^2=-1 so now you have 24+42i-27(-1) combine like terms to give you 51+42i And this is in standard form. :)